# CafkNet: GNN-Empowered Forward Kinematic Modeling for Cable-Driven Parallel Robots

### 28 Feb 2024 v1

@@ Abstract
When deploying Cable-Driven Parallel Robots (CDPRs) in practice, one of the challenges is kinematic modeling. Unlike serial mechanisms, CDPRs have a simple inverse kinematics problem but a complex forward kinematics (FK) issue. Therefore, the development of accurate and efficient FK solvers has been a prominent research focus in CDPR applications. By observing the topology within CDPRs, in this letter, we propose a graph-based representation to model CDPRs and introduce CafkNet, a fast and general FK solver, leveraging Graph Neural Network (GNN). Extensive experiments are conducted on 3D and 2D CDPRs across various configurations, including under-constrained, fully-constrained, and over-constrained cases, in both simulation environments and real-world scenarios. The experimental results showcase that CafkNet can learn the internal topological information of CDPRs and accurately solve the FK problem as an FK solver. Furthermore, training the CafkNet model on partial configurations enables zero-shot generalization to other configurations. Lastly, CafkNet effectively bridges the sim2real gap by using both simulation data and part of real-world data. To the best of our knowledge, it is the first study that employs the GNN to solve FK problem for CDPRs.
@@ I Introduction
It is obvious that due to the nonlinearity and complexity of FK problems, FK solvers have a large computational burden and also suffer from the problem of no solution or multiple solutions. Furthermore, configuration-dependent FK solvers cannot be easily applied to the reconfiguration CDPR (RCDPR) [11] or the case of simply adding or subtracting some cables. However, according to observations in [12], we notice that the change of CDPR’s configuration does not affect the topological architecture if its configuration can be represented as a graph. In addition, it is an indisputable fact that the Sim2Real gap is inevitable, despite significant efforts to reduce differences [13, 14, 15]. This is due to the difficulty of measuring and modeling systematic errors (cable attachment position errors) and random errors (cable sagging) in CDPR, some of which are even unknowable. One promising approach currently being explored is using real data to learn about these errors, as it has been shown to be an effective way to bridge the gap.
In this paper, we propose an FK solver based on Graph Neural Networks (GNN) for CDPR, denoted by CafkNet. By utilizing the proposed graph representation of CDPR, the CafkNet can rapidly learn the underlying topological structure of CDPR and serve as an FK solver to solve FK problems. Thanks to the topological similarity of graphs, CafkNet trained from a specific configuration can be easily transferred to other configurations without fine-tuning. Additionally, CafkNet performs well in bridging the Sim2Real gap between vast easily accessible simulation data and limited expensive real-world data. We have conducted extensive simulations and real-world experiments to demonstrate the modeling capabilities of CafkNet. To the best of our knowledge, it is the first study to employ the GNN for FK problem of CDPRs.
@@
  • We propose a forward kinematic modeling method based on GNN, CafkNet, which is the first GNN-empowered FK solver for CDPRs.
  • We use graph similarity to enable zero-shot and few-shot transfer learning, allowing CafkNet to quickly transfer its knowledge from one configuration to others.
  • We employ CafkNet to bridge the Sim2Real gap in CDPR, which has been verified in simulation and real machine experiments.
The rest of the paper is organized as follows. Sec. II reviews related studies about KM of CDPR and GNN applications in robotics. Sec. III details the CDPR geometric model. Sec. IV introduces the proposed graph representation of CDPR and the architecture of CafkNet. Then, Sec. V demonstrates the experiment details followed by modeling results of CafkNet in Sec. VI. Finally, the conclusion and outlook are presented in Sec. VII.
Figure 1: Workflow of CafkNet. (a) The geometric model of a CDPR with $m$ cables. (b) The proposed graph presentation for the CDPR.
@@ IV Methodology
IV-A Graph Representation of CDPRs
A graph typically consists of nodes and edges. In this work, we define three types of nodes and two types of edges to represent the CDPR to a graph, as outlined in Fig. 1-(b). Based on the geometric structure of CDPR in Fig. 1-(a), the framework can be considered as a world node $\mathbf{N}_{w}$, and the end-effector is seen as a body node $\mathbf{N}_{b}$. Furthermore, the position of the world frame $\{O\}$ attached to the framework, i.e., $[0,0,0]$, is set as the initial feature of $\mathbf{N}_{w}$. Similarly, the pose $\mathbf{Q}$ of the body frame attached to the end-effector is selected as the node feature of $\mathbf{N}_{b}$. For forward kinematic modeling, $\mathbf{Q}$ is unknown at first, so its node feature is initialized to be $[0,0,0,0,0,0]$. For CDPRs, cables connecting the framework and the end-effector are defined as a set of cable nodes $\mathbf{N}_{c}=\{N_{c}^{1},N_{c}^{2},\cdots,N_{c}^{m}\}$ for $m$ cables. Here $N_{c}^{i}$ refers to the $i$-th cable node, and its initial feature is the cable length, i.e., the scalar $||\mathbf{l}_{i}||$ in (1). In addition, the connection of the cable to framework and the connection to end-effector are considered as two types of edges, $\mathbf{E}_{wc}$ and $\mathbf{E}_{cb}$, whose edge features $e^{i}_{wc}$ and $e^{i}_{cb}$ for $i$-th cable are the attachment position vectors $A_{i}$ and $B_{i}$ in (1), respectively. As such, the CDPR can be formulated as a graph $G$ as
@@ IV-B GNN-Empowered Forward Kinematic Modeling
In this subsection, we will parameterize the FK problem of CDPRs as a GNN using the Encoder-Propagation-Decoder format [30] for message passing, as summarised in Algo. 1.
@@
Message Propagation: The propagation model consists of $N$ stacked message propagation blocks to update node embedding, which means message computation and aggregation are recurrently applied $N$ times. In each message propagation block, a top-down message-passing scheme is adopted. Specifically, the model first computes the message from the world node to each cable node $i$ by
@@
where $\gamma_{cb}$ denotes the MLP from cable to body nodes. After messages of all cable nodes $\{m^{i}_{c,b}\}$ are calculated, the body node is updated via
@@ V-A Dataset
In this work, we take $8$ configurations of CDPR into consideration, including spatial ones (e.g., Fig. 2-(a) and (b)) and planar ones (e.g., Fig. 2-(c) and (d)). We investigate under-constrained, fully-constrained and over-constrained cases, ranging from four cables to ten cables. The cable attachment locations in each configuration are detailed in Tab. I, where $A_{i}$ refers to $\mathbf{p}_{A_{i}}$ in (1), and $B_{i}$ indicates $\mathbf{v}_{i}$ in (1). In addition, we validate our method in the real experiment setup, as shown in Fig. 2-(c)
@@
TABLE I: Cable Attachment Locations for Various Configurations
Config. SimC4 SimC5 SimC6 SimC7 SimC8 SimC9 SimC10 ExpC4
$A_{i}$ (0.0, 1.0, 1.0) (0.5, 0.5, 1.0) (0.5, 1.0, 0.0) (1.0, 1.0, 1.0) (0.1, 1.0, 0.0) (1.0, 1.0, 1.0) (0.0, 1.0, 0.0) (0.0, 1.0, 1.0) (0.1, 1.0, 0.0) (1.0, 0.0, 1.0) (0.1, 1.0, 0.0) (1.0, 1.0, 1.0) (0.0, 0.0, 0.0)
(1.0, 1.0, 1.0) (0.0, 1.0, 1.0) (0.5, 0.0, 0.0) (1.0, 0.0, 1.0) (0.9, 1.0, 0.0) (1.0, 0.0, 1.0) (1.0, 1.0, 0.0) (1.0, 1.0, 1.0) (0.9, 1.0, 0.0) (0.0, 0.0, 1.0) (0.9, 1.0, 0.0) (1.0, 0.0, 1.0) (1.0, 0.0, 0.0)
(m) (1.0, 0.0, 1.0) (1.0, 1.0, 1.0) (0.0, 1.0, 1.0) (0.0, 0.0, 1.0) (0.5, 0.0, 0.0) (0.0, 0.0, 1.0) (1.0, 0.0, 0.0) (1.0, 0.0, 1.0) (0.5, 0.0, 0.0) (0.0, 0.5, 0.0) (0.9, 0.0, 0.0) (0.0, 0.0, 1.0) (1.0, 1.0, 0.0)
(0.0, 0.0, 1.0) (1.0, 0.0, 1.0) (0.0, 1.0, 1.0) (0.0, 0.0, 0.0) (0.0, 0.0, 1.0) (0.0, 1.0, 1.0) (1.0, 0.5, 0.0) (0.1, 0.0, 0.0) (0.0, 0.5, 0.0) (0.0, 1.0, 0.0)
(0.0, 0.0, 1.0) (1.0, 1.0, 1.0) (0.0, 1.0, 1.0) (1.0, 0.5, 0.0)
$B_{i}$ (-30, 30, 15) (0, 0, 15) (0, 30, 15) (30, 30, -15) (-25.5, 30, 15) (30, 30, -15) (-30, 30, -15) (-30, 30, 15) (-25.5, 30, 15) (30, -30, -15) (-25.5, 30, 15) (30, 30, -15) (150, -100, 0)
(30, 30, 15) (-30, 30, 15) (0, -30, 15) (30, -30, -15) (25.5, 30, 15) (30, -30, -15) (30, 30, -15) (30, 30, 15) (25.5, 30, 15) (-30, -30, -15) (25.5, 30, 15) (30, -30, -15) (-150, -100, 0)
(mm) (30, -30, 15) (30, 30, 15) (-30, 30, -15) (-30, -30, -15) (0, -30, 15) (-30, -30, -15) (30, -30, -15) (30, -30, 15) (0, -30, 15) (-30, 0, 15) (25.5, -30, 15) (-30, -30, -15) (-150, 100, 0)
(-30, -30, 15) (30, -30, 15) (-30, 30, -15) (-30, -30, -15) (-30, -30, 15) (-30, 30, -15) (30, 0, 15) (-25.5, -30, 15) (-30, 0, 15) (150, 100, 0)
(-30, -30, 15) (30, 30, -15) (-30, 30, -15) (30, 0, 15)
V-B Implementation Details
In this work, we use a supervised learning paradigm to train CafkNet for CDPRs based on the above dataset. Here, the Root Mean Square Error (RMSE) is selected as our performance metric. Note that we only calculate the RMSE loss about positions, disregarding any variations in orientations. In addition, Adam [31] is utilized as the optimizer, with a batch size of $32$. Other parameters, such as the hidden feature dimensions and number of MLP layers in CafkNet, the initial learning rate, and the reduction factor for the learning rate scheduler, are optimized automatically via the toolkit in [32]. All codes are developed using PyTorch and trained on a single NVIDIA GeForce RTX 4090 GPU card.
@@ VI-A FK Modeling for CDPR
Figure 3: Reference trajectories and resulting trajectories solved by the MLP and CafkNet (including one2one and multi-task cases) at configurations SimC4 to SimC10. Top row: MLP; Middle row: CafkNet (one2one); Bottom row: CafkNet (multi-task).
The results of modeling accuracy, i.e., the RMSE loss of predicted positions and reference positions, as revealed in Tab. II. Here we can observe that the opt. FK solver achieves the smallest error in SimC6 to SimC10 but fails in SimC4 and SimC5, where CafkNet achieves the lowest error. But from SimC6 to SimC10, the errors of CafkNet remain within $1.5$ mm. On the other hand, the MLP solver exhibits the poorest accuracy, with errors exceeding $15$ mm. Furthermore, we notice that the accuracy of CafkNet (multi-task) is comparable to CafkNet (one2one).
However, considering the computational time for FK solving, as shown in Tab. III, the opt. FK solver takes an average time of approximately $1$ s for the best-accuracy configurations, SimC4 $\sim$ SimC10, but experiences a significant time raising at failure cases. In contrast, it can be found that CafkNet (one2one) achieves remarkably low computation time, with only $0.649$ s for SimC4 and a few milliseconds for other configurations. Considering both computational time and solution accuracy, the CafkNet offers a trade-off between low time consumption and relatively high precision, demonstrating significant potential for real-time CDPR applications compared to the opt. FK solver.
Note that the Tab. III indicates the overall time taken by both methods to compute a complete trajectory, comprising 100 FK problems. The resulting trajectories are visually presented in Fig. 3, illustrating the FK outcomes of MLP (top row), CafkNet (one2one) (middle row), and CafkNet (multi-task) (bottom row). The close correspondence observed between the predicted and reference trajectories serves as evidence of the accurate FK solutions.
TABLE II: Modeling Accuracy (unit: mm).
@@
TABLE III: Time Cost of FK solving (unit: s).
@@ VI-B Transfer Learning in CDPR
In this subsection, we investigate the transferability of our CafkNet, and the resulting modeling errors are given in Tab. IV. In the GNN, all cable and framework node interactions share parameters $\sigma_{C}$, which allows us to transfer the kinematic model to CDPRs with different cable configurations to assess the generalization ability of the trained model. We conduct a series of experiments on transferability tests, where the kinematic model is transferred without further fine-tuning.
@@
For example, the second row in Tab. IV represents the training of CafkNet on the data from SimC4, followed by directly solving the FK problems on SimC5-SimC10, respectively. It can be observed that the errors in Exp C-1 are relatively high. This result aligns with our expectations, as CafkNets have limited and singular learning data on a single configuration, resulting in better performance on the same configuration and poorer performance on other configurations.
@@
TABLE IV: Modeling Accuracy on Transferability of CafkNet (unit: mm). Bold: minimum test error in each configuration.
@@ VI-C1 Simulated Data with Artificial Noises
In Fig. 2-(c), we find that the pose of CDPR can be accurately captured by the motion capture system, so we do not add noise to the pose in the simulated data. On the contrary, the main errors in the real experiment come from the length of the cable, including the error of the motor encoder, as well as the position of the cable on the pulley and where it is mounted at the end-effector. Thus, we introduce the random error $\epsilon$ into the simulator mainly from the position noises in attachment positions, i.e., $A_{i}\pm\epsilon,~{}B_{i}\pm\epsilon$, where $\epsilon\sim N(\mu=0,\sigma^{2})$. We conduct two sets of experiments targeting different $\sigma$ values to assess the performance of CafkNet when exposed to artificial noise, i.e., $\sigma=5$ and $\sigma=10$, respectively. In each experimental set, we also consider three categories based on the source of the training and testing data for each CDPR configuration. These categories include:

### 05 Mar 2024 v2

@@ Abstract
The Cable-Driven Parallel Robots (CDPRs) have gained significant attention due to their high payload capacity and large workspace. When deploying CDPRs in practice, one of the challenges is kinematic modeling. Unlike serial mechanisms, CDPRs have a simple inverse kinematics problem but a complex forward kinematics (FK) issue. Therefore, the development of accurate and efficient FK solvers has been a prominent research focus in CDPR applications. By observing the topology within CDPRs, in this paper, we propose a graph-based representation to model CDPRs and introduce CafkNet, a fast and general FK solver, leveraging Graph Neural Network (GNN). CafkNet is extensively tested on 3D and 2D CDPRs in different configurations, both in simulators and real scenarios. The results demonstrate its ability to learn CDPRs’ internal topology and accurately solve the FK problem. Then, the zero-shot generalization from one configuration to another is validated. Also, the sim2real gap can be bridged by CafkNet using both simulation and real-world data. To the best of our knowledge, it is the first study that employs the GNN to solve FK problem for CDPRs. Videos and codes are available at https://sites.google.com/view/cafknet/site.
@@ I Introduction
It is obvious that due to the nonlinearity and complexity of FK problems, FK solvers have a large computational burden and also suffer from the problem of no solution or multiple solutions. Furthermore, configuration-dependent FK solvers cannot be easily applied to the reconfiguration CDPR (RCDPR) [11] or the case of simply adding or subtracting some cables. However, according to our observations in [12], we notice that the change of CDPR’s configuration does not affect the topological architecture if its configuration can be represented as a graph. In addition, it is an indisputable fact that the Sim2Real gap is inevitable, despite significant efforts to reduce differences [13, 14, 15]. This is due to the difficulty of measuring and modeling systematic errors (e.g., cable attachment position errors) and random errors (e.g., cable sagging) in CDPR, some of which are even unknowable. One promising approach currently being explored is using real data to learn about these errors, as it has been shown to be an effective way to bridge the gap.
In this paper, we propose a forward kinematics solver based on graph neural Networks (GNN) for Cable-driven parallel robots, denoted by CafkNet. By utilizing the proposed graph representation of CDPR, the CafkNet can rapidly learn the underlying topological structure of CDPR and serve as an FK solver to solve FK problems. Thanks to the topological similarity of graphs, CafkNet trained from a specific configuration can be easily transferred to other configurations without fine-tuning. Additionally, CafkNet performs well in bridging the Sim2Real gap between vast easily accessible simulation data and limited expensive real-world data. We have conducted extensive simulations and real-world experiments to demonstrate the modeling capabilities of CafkNet. To the best of our knowledge, it is the first study to employ the GNN for FK problem of CDPRs.
@@
  • We propose the CafkNet, a fast and general FK solver for CDPRs with any configurations based on GNN.
  • We use graph similarity to enable zero/few-shot transfer learning, allowing CafkNet to quickly transfer learned FK knowledge from one configuration to another.
  • We employ CafkNet to bridge the Sim2Real gap in CDPRs, which has been verified in simulation and real machine experiments.
The overview of CafkNet is exhibited in Fig. 1. The rest of the paper is organized as follows. Sec. II reviews related studies about KM of CDPR and GNN applications in robotics. Sec. III details the CDPR geometric model. Sec. IV introduces the proposed graph representation of CDPR and the architecture of CafkNet. Then, Sec. V demonstrates the experiment details followed by modeling results of CafkNet in Sec. VI. Finally, the conclusion and outlook are presented in Sec. VII.
Figure 1: Overview of CafkNet. (a) The geometric model of a spatial CDPR with $m$ cables. (b) The proposed graph presentation for the CDPR (see Sec. IV-A). (c) The top-down computation scheme in each message propagation block. The cable nodes collect data from the world node and are eventually aggregated by the body node (see Sec. IV-B).
@@ IV Methodology
IV-A Graph Representation of CDPR
A graph typically consists of nodes and edges. In this work, we define three types of nodes and two types of edges to represent the CDPR as a graph, as outlined in Fig. 1-(b). Based on the geometric structure of CDPR in Fig. 1-(a), the framework can be considered as a world node $\mathbf{N}_{w}$, and the end-effector is seen as a body node $\mathbf{N}_{b}$. Furthermore, the position of the world frame $\{O\}$ attached to the framework, i.e., $[0,0,0]$, is set as the initial feature of $\mathbf{N}_{w}$. Similarly, the pose $\mathbf{Q}$ of the body frame attached to the end-effector is selected as the node feature of $\mathbf{N}_{b}$. For FK modeling, $\mathbf{Q}$ is unknown, so its node feature is initialized to be $[0,0,0,0,0,0]$. For CDPRs, cables connecting the framework and the end-effector are defined as a set of cable nodes $\mathbf{N}_{c}=\{N_{c}^{1},N_{c}^{2},\cdots,N_{c}^{m}\}$ for $m$ cables. Here $N_{c}^{i}$ refers to the $i$-th cable node, and its initial feature is the cable length, i.e., the scalar $||\mathbf{l}_{i}||$ in (1), which is known in FK modeling. In addition, the connection of the cable to framework and the connection to end-effector are considered as two types of edges, $\mathbf{E}_{wc}$ and $\mathbf{E}_{cb}$, whose edge features $e^{i}_{wc}$ and $e^{i}_{cb}$ for $i$-th cable are the attachment position vectors $A_{i}$ and $B_{i}$ in (1), respectively. As such, the CDPR can be formulated as a graph $G$ as
@@ IV-B GNN-Empowered Forward Kinematic Modeling
In this subsection, we will parameterize the FK problem of CDPRs as a GNN using the Encoder-Propagation-Decoder format [30] for message passing, as summarized in Algo. 1.
@@
Message Propagation: The propagation model consists of $N$ stacked message propagation blocks to update node embedding, which means message computation and aggregation are recurrently applied $N$ times. In each message propagation block, a top-down message-passing scheme is adopted, as exhibited in Fig. 1-(c). Specifically, the model first computes the message from the world node to each cable node $i$ by
@@
where $\gamma_{cb}$ denotes the MLP from cable to body nodes, and $h_{b}$ are hidden features of the body node. Then $e^{i}_{cb}$ refers to the edge feature of edge from $i$-th cable node to body node. After messages of all cable nodes $\{m^{i}_{c,b}\}$ are calculated, the body node is updated via
@@ V-A Dataset
In this work, we consider $8$ configurations of CDPR, including spatial ones (e.g., Fig. 2-(a) and (b)) and planar ones (e.g., Fig. 2-(c) and (d)). We investigate under-constrained, fully-constrained and over-constrained cases, ranging from four cables to ten cables. In addition, we validate our method in the real experiment setup, as shown in Fig. 2-(c)
@@
V-B Metrics
In this work, we use a supervised learning paradigm to train CafkNet for CDPRs based on the above dataset. Here, the Root Mean Square Error (RMSE) is selected as our performance metric. Note that we only calculate the RMSE loss about positions, disregarding any variations in orientations.
V-C Implementation Details
In our implementation, we utilize Adam [31] is utilized as the optimizer, with a batch size of $32$. Other parameters, such as the hidden feature dimensions and number of MLP layers in CafkNet, the initial learning rate, and the reduction factor for the learning rate scheduler, are optimized automatically via the toolkit in [32]. All codes are developed using PyTorch and trained on a single NVIDIA GeForce RTX 4090 GPU card.
@@ VI-A FK Modeling for CDPR
The results of modeling accuracy, i.e., the RMSE loss of predicted positions and reference positions, as revealed in Tab. I. Here we can observe that the opt. FK solver achieves the smallest error in SimC6 to SimC10 but fails in SimC4 and SimC5, where CafkNet achieves the lowest error. But from SimC6 to SimC10, the errors of CafkNet remain within $1.5$ mm. On the other hand, the MLP solver exhibits the poorest accuracy, with errors exceeding $15$ mm. Furthermore, we notice that the accuracy of CafkNet (multi-task) is comparable to CafkNet (one2one).
However, considering the computational time for FK solving, as shown in Tab. II, the opt. FK solver takes an average time of approximately $1$ s for the best-accuracy configurations, SimC4 $\sim$ SimC10, but experiences a significant time raising at failure cases. In contrast, it can be found that CafkNet (one2one) achieves remarkably low computation time, with only $0.649$ s for SimC4 and a few milliseconds for other configurations. Considering both computational time and solution accuracy, the CafkNet offers a trade-off between low time consumption and relatively high precision, demonstrating significant potential for real-time CDPR applications compared to the opt. FK solver.
Note that the Tab. II indicates the overall time taken by both methods to compute a complete trajectory, comprising 100 FK problems. The resulting trajectories are visually presented in Fig. 3, illustrating FK solutions solved by MLP, CafkNet (one2one), and CafkNet (multi-task), respectively. The close correspondence observed between the predicted and reference trajectories serves as evidence of the accurate FK solutions.
TABLE I: Modeling Accuracy (unit: mm).
@@
TABLE II: Time Cost of FK solving (unit: s).
@@
Figure 3: Reference trajectories and resulting trajectories solved by MLP and CafkNet under configurations (a) - (g). Top row: by MLP; Middle row: by CafkNet (one2one); Bottom row: by CafkNet (multi-task).
@@ VI-B Transfer Learning in CDPR
In this subsection, we investigate the transferability of our CafkNet, and the resulting modeling errors are given in Tab. III. In the GNN, all cable and framework node interactions share parameters $\sigma_{C}$, which allows us to transfer the kinematic model to CDPRs with different cable configurations to assess the generalization ability of the trained model. We conduct a series of experiments on transferability tests, where the kinematic model is transferred without further fine-tuning.
@@
For example, the second row in Tab. III represents the training of CafkNet on the data from SimC4, followed by directly solving the FK problems on SimC5-SimC10, respectively. It can be observed that the errors in Exp C-1 are relatively high. This result aligns with our expectations, as CafkNets have limited and singular learning data on a single configuration, resulting in better performance on the same configuration and poorer performance on other configurations.
@@
TABLE III: Modeling Accuracy on Transferability of CafkNet (unit: mm). Bold: minimum test error in each configuration.
@@ VI-C1 Simulated Data with Artificial Noises
In Fig. 2-(c), we find that the pose of CDPR can be accurately captured by the motion capture system, so we do not add noise to the pose in the simulated data. On the contrary, the main errors in the real experiment come from the length of the cable, including the error of the motor encoder, as well as the position of the cable on the pulley and where it is mounted at the end-effector. Thus, we introduce the random error $\epsilon$ in the simulator mimicking cable attachment position noises, i.e., $A_{i}\pm\epsilon,~{}B_{i}\pm\epsilon$, where $\epsilon\sim\mathcal{N}(\mu=0,\sigma^{2})$ and $\mathcal{N}$ refers to the Gaussian distribution. We conduct two sets of experiments targeting different $\sigma$ values to assess the performance of CafkNet when exposed to artificial noise, i.e., $\sigma=5$ and $\sigma=10$, respectively. In each experimental set, we also consider three categories based on the source of the training and testing data for each CDPR configuration. These categories include:
@@
Figure 4: Modeling accuracy of CafkNet working on artificial noise data. Here errors in case 1 to case 3 are generated by $\epsilon\sim\mathcal{N}(0,5^{2})$, and errors in case 4 to case 6 are generated by $\epsilon\sim\mathcal{N}(0,10^{2})$. $\mathcal{N}$ refers to the Gaussian distribution.
Figure 5: Data sources of Sim2Real experiment for the configuration of ExpC4. (i) Simulated data. In the simulator, there are $100$ random trajectories and $100$ samplings per trajectory. (ii) Clean data. In the simulator, we have devised a target trajectory (dashed curve) to control the real CDPR to execute it. Here $1000$ samplings are generated. (iii) Noise data. In the real setup, the end-effector of CDPR follows the target trajectory. During the experiment, a total of 1103 data are collected. Above all, each data composes the end-effector pose ($x,y,\theta$) and the corresponding length of $4$ cables.